OpenStudy (firejay5):

The initial # of bacteria in a culture is 12,000. The # after 4 hours of the application of a new drug is 4,000. 1. Write an exponential function to model the population y of bacteria after x hours. 2. How many bacteria are there after 24 hours?

OpenStudy (nincompoop):

would you like to call for a QH help?

OpenStudy (anonymous):

p=P x e^(kt) where p=population, P=initial population, k=constant, t=time 1. y=12000e^(kt) 2. 4000=12000e^4k e^4k=4000/12000=1/3 4k=ln(1/3) therefore k=[ln(1/3)]/4 To find how many bacteria after 24hours. p=12000e^{24x[ln(1/3)]/4} and you'll get 16.46 population

OpenStudy (firejay5):

@SherBear0101 Exponential function: y = A * B^X

OpenStudy (anonymous):

yes, what about it?

OpenStudy (firejay5):

@SherBear0101 Is the work that you did works for y = A * B^x, could you work it out using the function I gave you.

OpenStudy (anonymous):

it should work the same just by replacing the letters